Suppose that the price elasticity of demand for toffee is 1.75. A drop in price will result in:

A. an increase in total toffee revenue

B. a decrease in the quantity of toffee

C. no change in total toffee revenue

D. a decrease in total toffee revenue

E. None of the options
Answer #1

Elasticity is the ratio of percentage change in QUANTITY demanded and percentage change in price.

Total revenue multiplication of price and quantity.

Elasticity of 1.75 ,means 1% Decrease in demand , Increase QUANTITY by 1.75 %.

TR=P*Q

%∆ TR=%∆p+%∆Q

%∆P=-1%

%∆Q=+1.75%

%∆TR=-1+1.75=0.75%

So 1% Decrease in price will increase total revenue by 0.75%.

Option A is correct